A muon formed high in the Earth’s atmosphere is measured by an observer on the Earth’s surface to travel at speed 𝑣 = 0.990𝑐 for a distance of 4.60 km before it decays into an electron, a neutrino, and an antineutrino (𝜇− → 𝑒− + 𝜈 + 𝜈̅).  

(a) For what time interval does the muon live as measured in its reference frame?  

(b) How far does the Earth travel as measured in the frame of the muon? 


1
Expert's answer
2021-11-16T10:14:06-0500

"\text {Lorentz factor:}"

"v = 0.99c"

"\gamma=\frac{1}{\sqrt{1-\frac{v^2}{c^2}}}=\frac{1}{\sqrt{1-0.99^2}}\approx7.09"

"\text{for Earthbound observer:}"

"s_E = 4.6km"

"v_E= 0.99c"

"c \approx 3*10^8\frac{m}{c}"

"t_E= \frac{s_E}{v_E}= \frac{4600}{0.99*3*10^8}\approx1.54*10^{-5}\ s"

"\text{for muonbound observer:}"

"t_m = \frac{t_E}{\gamma}=\frac{1.54*10^{-5}}{7.09}\approx 2.18*10^{-6}s"

"s_m = vt_m= v* \frac{t_E}{\gamma}=\frac{s_E}{\gamma}\approx648.8m"

"\text{Answer:}"

"a)1.54*10^{-5}\ s"

"b)648.8\ m"